OpenStudy (anonymous):

********MEDALS********* 3x-4=2x+7

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

3x-2x=7+4 x=11

OpenStudy (anonymous):

3x-4=2x+7 3x-2x = 7+4 x = 11

OpenStudy (anonymous):

this is a baby qustion.

OpenStudy (anonymous):

and that's why im asking u guys because I don't want to do it HEY, I'M OUT OF SCHOOL XD

OpenStudy (anonymous):

@felavin did she ever said that she's doing phD in maths :P

OpenStudy (anonymous):

XD hitman

OpenStudy (anonymous):

just a need a little thinking not ever ask a question

OpenStudy (anonymous):

well did u want a medal ? if u didn't want to answer it then don't XD